Given the following system of equations
x+y=5k
2x=4k
where k is a constant real number, solve the system for (x,y)in terms of k.
Do not replace the constant k with a numerical value.
Give your solution as an ordered pair in terms of k.









Answer by rfer(16322)

2x=4k
x=2k
--------------
x+y=5k
2k+y=5k
y=3k
---------------
x+3k=5k
x=2k
----------
(2k,3k)